Jenny is planning to deposit ₱17,000.00 Quezon Metropolitan Bank is offering 7.5% compounded semi-annually

for 5 years while Quezon Premier Bank is offering 7% compounded monthly for 5 years. Which bank should she

deposit her money? Justify your answer by computing and comparing the maturity value for each bank.


1
Expert's answer
2022-02-10T16:51:01-0500

"FV=PV(1+\frac{r}{n})^{n×t}"


For Quezon Metropolitan Bank:

PV= ₱17,000.00

r = 7.5%

t = 5years

n= 2


"FV = 17,000.00(1+""\frac{0.075}{2})^{2×5}"


"=17,000.00(1.0375)^{10}=\u20b1 24565.75"

For Quezon Premier Bank:

PV= ₱17,000.00

r = 7%

t = 5years

n= 12


"FV = 17,000.00(1+""\frac{0.07}{12})^{12×5}"


"=17,000.00(1.005833)^{60}=\u20b1 24099.15"

She should deposit her money with Quezon Metropolitan Bank since it has higher maturity value compared to Quezon Premier Bank